1/3 (6ab-5ab)-1/6(2b-12)
1
Expert's answer
2013-04-17T08:05:13-0400
we know that for every x, a,b ax+bx=(a+b)x then (6ab-5ab)=(6-5)ab =1*ab=abalso (2b-12)=(b-6)*2

hence 1/3 (6ab-5ab)=1/3 *ab=ab/31/6(2b-12)=2(b-6)/6=(b-6)/3
so we obtain
1/3 (6ab-5ab)-1/6(2b-12)=ab/3 -(b-6)/3
(b-6)/3=b/3-6/3=b/3-2
ab/3 -(b-6)/3=ab/3 -(b/3 -2)
-(b/3-2)=-b/3+2
Therefore, ab/3 -(b/3-2)=ab/3-b/3+2=(a-1)*b/3 +2 The result is (a-1)*b/3+2